For a certain ideal gas, R = 0.200 KJ/kg-K and k = 1.384. What are the values of Cp and Cv when added in KJ/kg-K?


1
Expert's answer
2021-06-08T09:34:08-0400

For an ideal gas


"C_p-C_V=R=0.200""\frac{C_p}{C_V}=k=1.384"

Roots


"C_p=0.721\:{\rm KJ/kg\cdot K}, \quad C_V=0.521\:\rm KJ/kg\cdot K"